# An Excited Dark Matter Solution to the MeV Galactic Center Excesses

### v1

@@ 1 Introduction
The Galactic 511 keV annihilation line has remained one of the most persistent puzzles in high-energy astrophysics for decades. INTEGRAL/SPI measurements (Prantzos:2010wi; kierans2019positron; Siegert:2019tus; Siegert_2023) reveal a bright, nearly symmetric emission around the Galactic Center (GC) and extending $\sim 6^{\circ}$$8^{\circ}$ on the sky. The inferred positron injection rate, $R_{e^{+}}\!\sim\!10^{43}\,{\rm s^{-1}}$ (Prantzos:2010wi), together with the narrow extension of the signal, implies injection energies below a few–tens of MeV.
@@
While $\beta^{+}$ decays of 26Al, 44Ti and 56Ni can account for a substantial fraction of the disk component (Diehl:2005py; Prantzos:2010wi), conventional astrophysical populations struggle to reproduce the bright, compact bulge emission (Siegert_2023). Stellar templates such as the nuclear stellar bulge or a boxy bulge can be tuned to match the observed morphology (Siegert:2021trw), which has been interpreted as support for a stellar origin. However, magnetohydrodynamic simulations of Milky Way–like galaxies show that dark matter (DM)–baryon interactions can drive the DM distribution into a configuration that closely traces the stellar bulge and boxy bulge structure (Silk). In such cases, template fits based solely on stellar tracers cannot reliably distinguish a purely stellar explanation from one in which DM contributes significantly.
@@
The DM interpretation of the 511 keV line has therefore received sustained attention (Boehm:2003bt; Vincent:2012an; DelaTorreLuque:2024fcc; DelaTorreLuque:2023cef; laTorreLuquePedro:2024est). Simple $s$-wave annihilation scenarios, in which MeV-scale DM directly produces $e^{+}e^{-}$ pairs, are tightly constrained by CMB and Big Bang Nucleosynthesis (Beacom:2005qv; Sizun:2007ds; Wilkinson:2016gsy; DelaTorreLuque:2024fcc), disfavoring standard thermal MeV DM as the source of the positrons.
@@
A particularly promising alternative is the Excited Dark Matter (XDM) framework proposed in Finkbeiner:2007kk. In this scenario, collisions between DM particles inelastically excite one particle to a slightly heavier state that decays back via a light mediator, producing an $e^{+}e^{-}$ pair. Because excitation requires a threshold in relative velocity, the positron production rate acquires an exponential dependence on the local velocity dispersion. As the halo velocity dispersion decreases with galactocentric radius, the upscattering rate is strongly suppressed outside the inner kiloparsec, yielding a bulge-peaked morphology that naturally declines toward the disk. Quantitative studies by Vincent:2012an and Cappiello:2023qwl showed that such XDM models can reproduce the INTEGRAL/SPI angular profile for TeV-scale DM masses, providing a markedly better fit than standard annihilating DM scenarios.
@@
In parallel, a reanalysis of archival COMPTEL data has uncovered a broad excess around 2–3 MeV, detected at $\sim 8\sigma$ significance and consistent with in-flight annihilation (IfA) of $\sim 2$ MeV positrons (Knodlseder:2025pnx). The spatial distribution of this MeV excess closely tracks the 511 keV emission, strongly suggesting a common source of low-energy positrons.
@@
A third anomaly arises in the Central Molecular Zone (CMZ), the dense neutral-gas complex within the central few hundred parsecs of the Galaxy. Multiple molecular tracers indicate an H2 ionization rate $\zeta_{\mathrm{H}_{2}}\gtrsim 10^{-15}\,\mathrm{s^{-1}}$ (Oka_2005; Oka_2020), which is difficult to reconcile with standard CRs (cosmic rays): they are strongly attenuated in dense clouds and their predicted ionization rate falls short by at least two orders of magnitude (Ravikularaman_2025; Huang_2021). Moreover, ionization powered by a single central source such as Sgr A* or HESS J1745-290 is expected to peak sharply near the source and decline steeply with radius, whereas the data favour a much flatter radial profile across the CMZ (Dogiel_2013; Ravikularaman_2025). DelaTorreLuque:2024fcc showed that MeV DM annihilating into $e^{\pm}$ can efficiently ionize the CMZ and reproduce both the level and approximate flatness of the ionization profile, in a region of parameter space compatible with existing constraints and with positron injection rates comparable to those required for the 511 keV line.
@@
The remainder of this paper is organised as follows. Section 2 summarizes the XDM mechanism, introduces the key phenomenological parameters and describes our implementation of the excitation rate and positron source term. Section 3 presents the fit to the 511 keV line, and the prediction of the IfA signal and the CMZ ionization contribution. Section 4 discusses the implications of our results and the prospects for testing this scenario with future MeV-band observations.
2 Excited Dark Matter Framework
@@
We consider a minimal dark sector with two nearly degenerate states: a stable ground state $\chi$ and a slightly heavier excited state $\chi^{*}$, separated by a small mass splitting $\Delta m=m_{\chi^{*}}-m_{\chi}$. This setup, first proposed for the 511 keV line in Finkbeiner:2007kk and revisited in Vincent:2012an; Cappiello:2023qwl, provides a simple way for DM to produce low-energy positrons in the Milky Way.
Collisions between DM particles can inelastically upscatter one of them to the excited state, $\chi\chi\rightarrow\chi\chi^{*}$ provided their relative velocity exceeds a kinematic threshold. The excited state then decays $\chi^{*}\rightarrow\chi e^{+}e^{-}$, releasing an $e^{+}e^{-}$ pair with total energy $\sim\Delta m$. In contrast to standard annihilating DM, this mechanism converts a small fraction of the halo kinetic energy into electron–positron pairs, leaving the overall DM abundance and large-scale halo structure unchanged. Even if only a tiny fraction of particles are excited per Hubble time, the kinetic energy reservoir is sufficient to power the observed positron production rate in the Galactic bulge (Prantzos:2010wi; Siegert_2023).
Each upscattering event injects positrons with kinetic energy $E_{e^{+}}\simeq\Delta m/2$, so $\Delta m\sim\mathrm{few\penalty 10000\ MeV}$ naturally yields non-relativistic positrons consistent with the narrow 511 keV line. The strong velocity threshold induces a steep radial dependence in the excitation rate: because the velocity dispersion $\sigma_{v}(r)$ decreases with galactocentric radius, the upscattering rate is exponentially suppressed beyond the inner kiloparsec. This produces a sharply peaked, bulge-dominated morphology that matches the $\sim 6-8^{\circ}$ FWHM of the INTEGRAL/SPI signal (Weidenspointner:2007rs; Siegert:2019tus) without requiring ad hoc modifications of the DM density profile.
@@
as shown in Finkbeiner:2007kk; Vincent:2012an; Cappiello:2023qwl. The exponential factor arises from integrating the inelastic cross section over the local Maxwell–Boltzmann velocity distribution and naturally suppresses the signal outside the Galactic bulge.
At the microscopic level, we parameterize the inelastic upscattering cross section as (Cappiello:2023qwl)
@@
where $\sigma_{\textrm{mr}}$ sets the normalization of the inelastic scattering cross section in the moderately relativistic limit. The threshold $v_{\mathrm{th}}$ introduces a sharp kinematic cutoff, confining excitations to the high-velocity tail of the halo distribution and imprinting the exponential suppression with radius seen in Eq. (2). This form approximates the more general expressions obtained in light-mediator models (Finkbeiner:2007kk; Pospelov:2007xh; Arkani-Hamed:2008hhe), while capturing the astrophysically relevant dependence needed to model the Galactic signal.
In what follows, we fold this threshold behaviour into a full computation of the velocity-averaged excitation rate using the local velocity dispersion $\sigma_{v}(r)$ derived from the Jeans equation (Section 2). This quantity controls both the radial morphology of the 511 keV emission and the total positron injection rate relevant for the IfA and CMZ analyses.
@@
Velocity-averaged excitation rate. Following Cappiello:2023qwl, the velocity-averaged inelastic excitation cross section is obtained by integrating the microscopic cross section over the local DM velocity distribution, {widetext}
@@
where $\sigma_{v}(r)$ and $v_{\mathrm{esc}}(r)$ are the local DM velocity dispersion and escape velocity, and $N_{\mathrm{esc}}$ normalizes the truncated Maxwellian distribution. The threshold velocity $v_{\mathrm{th}}$ is defined in Eq. (1). [^1] In Eq. (5) of Cappiello:2023qwl, a cross section $\langle\sigma v\rangle_{\rm mr}\simeq\sigma_{\textrm{mr}}v_{\rm eff}$ was introduced as a constant normalization with an effective halo velocity $v_{\rm eff}\sim 10^{-3}c$. Here we instead derive $\langle\sigma v\rangle(r)$ by explicitly integrating over the full velocity distribution, which is appropriate for a geometric cross section $\sigma_{\rm mr}$ and ensures a self-consistent radial dependence.
@@
The velocity dispersion $\sigma_{v}(r)$ and escape velocity $v_{\mathrm{esc}}(r)$ entering Eq. (5) are shown in Fig. 1. They are computed by solving the isotropic Jeans equation under spherical symmetry, following Vincent:2012an; Cappiello:2023qwl. The total gravitational potential includes contributions from DM, bulge, disk and Sgr A*. We adopt the same baryonic potential models and parameter choices as Vincent:2012an, which reproduce Milky Way rotation-curve data and yield $\sigma_{v}\sim 200$$250\penalty 10000\ \mathrm{km\,s^{-1}}$ and $v_{\mathrm{esc}}\sim 600$$700\penalty 10000\ \mathrm{km\,s^{-1}}$ in the inner kiloparsec.
Figure 1: DM velocity profiles as a function of radial distance from the GC for the velocity dispersion $\sigma_{v}$ (blue) and escape velocity $v_{\mathrm{esc}}$ (red) for NFW (solid) and cNFW (dashed) DM density distributions. The cNFW profile has inner slope $\gamma=1.2$.
@@
The angular distribution of 511 keV photons is obtained by integrating the local emissivity of diffuse positrons along the line of sight. Unlike previous morphology studies that assumed instantaneous annihilation at the production site (Vincent:2012an; Cappiello:2023qwl), we propagate positrons with non-zero kinetic energy using the DRAGON2 code, following DelaTorreLuque:2023cef. This includes diffusion, reacceleration and energy losses, and yields a steady-state positron distribution before annihilation.
@@
where $\frac{{d\phi_{e}}}{{dE}}(E_{\mathrm{th}})$ is the propagated positron flux at thermal energy $E_{\mathrm{th}}$, $\sigma^{\mathrm{ps}}$ is the charge-exchange cross section for positronium formation with hydrogen (JeanP_2009), and $n_{e}$ is the ambient electron density. We take $n_{e}=1\penalty 10000\ \mathrm{cm^{-3}}$ in the Galactic plane and adopt an exponential decrease above and below the plane as in laTorreLuquePedro:2024est, combining a Nakanishi warm-gas component (Nakanishi_2003) and a Ferriere hot-gas component (Ferriere_1998).
@@
Relativistic positrons injected in the interstellar medium can annihilate with electrons before becoming thermal, producing a continuum of $\gamma$ rays up to the positron energy. IfA has long been recognized as a probe of MeV DM scenarios (Stecker; Beacom:2005qv; Sizun:2006uh), since high injection energies generate an IfA flux that can overshoot the observed diffuse MeV background (COMPTEL1994; Bouchet_2011).
As discussed in Sec. 1, the reanalysis of COMPTEL data by Knodlseder:2025pnx reveals a MeV excess consistent with IfA of $\sim 2$ MeV positrons sharing the 511 keV spatial template. In our XDM scenario, these positrons are produced in $\chi^{*}\to\chi e^{+}e^{-}$ decays with injection energy $E_{e^{+}}\simeq\Delta m/2$. For monoenergetic injection, the differential IfA $\gamma$-ray flux can be related to the 511 keV intensity (Beacom:2005qv; Beacom:2004pe):
@@
where $f_{\mathrm{Ps}}\simeq 0.97$ is the positronium fraction (JeanP_2009; Bouchet:2010dj) and $P^{\mathrm{IfA}}(E_{\gamma};E_{0},n_{e})$ is the photon spectrum per injected positron of initial energy $E_{0}=\Delta m/2$ in a medium of density $n_{e}$.
Here we compute the IfA signal following laTorreLuquePedro:2024est, using the steady-state positron spectrum obtained from DRAGON2:
@@
where the first integral over $E^{\prime}$ accounts for the propagated positron energy distribution, bounded between $E_{\textrm{max}}$ (the maximum steady-state energy) and $m_{e}/2$ (Bartels:2017dpb). The factor $(1/N_{\rm pos})\,dN_{\rm pos}/dE^{\prime}$ gives the fraction of positrons at energy $E^{\prime}$ that contribute to the 511 keV emission. The hydrogen density $n_{\textrm{H}}$ sets the number of target electrons, $P_{E^{\prime}\to E}$ is the probability for a positron to produce a photon before cooling from $E^{\prime}$ to $E$ (Beacom:2005qv), and $P=P_{E^{\prime}\to m_{e}}$ is the total probability to emit a photon before reaching thermal energies.
@@
As discussed in Sec. 1, multiple molecular tracers indicate an H2 ionization rate in the CMZ of $\zeta_{\mathrm{H}_{2}}\gtrsim 10^{-15}\,\mathrm{s^{-1}}$ with an approximately flat radial profile (Oka_2005; Oka_2020; Dogiel_2013; Ravikularaman_2025), which is difficult to explain with standard CRs alone (Huang_2021; Ravikularaman_2025). In light of the MeV–DM explanation proposed by DelaTorreLuque:2024fcc, we now ask whether the same XDM setup that fits the 511 keV morphology and IfA signal can also contribute appreciably to the CMZ ionisation.
We follow DelaTorreLuque:2024fcc to compute the H2 ionization rate $\zeta_{\textrm{H}_{2}}$ caused by the $e^{\pm}$ pairs emitted in XDM de-excitations inside the CMZ. The rate is
@@
where the prefactor 2 accounts for ionization by both $e^{-}$ and $e^{+}$, $E_{\textrm{max}}=\Delta m/2$ is the injection energy, and $E_{\textrm{min}}$ is the threshold for ionization in H2 ($15.43$ eV (Padovani_2009)). The cross section $\sigma(E)$ is the electron–H2 ionization cross section, calculated as in Eq. (7) of Padovani_2009, and $\theta_{e}(E)$ is the number of secondary ionizations per primary ionization, from Eq. (2.23) of krause2015crimecosmicray. The flux $J(E,\textbf{x})$ is the propagated $e^{\pm}$ flux from DRAGON2, which includes the strong energy losses suffered by these particles in the CMZ. Typical propagation distances before thermalization are of order tens of parsecs, depending on the injection energy.

### v2

@@ I Introduction
The Galactic 511 keV annihilation line has remained one of the most persistent puzzles in high-energy astrophysics for decades. INTEGRAL/SPI measurements (Prantzos and others, 2011; Kierans and others, 2019; Siegert et al., 2019; Siegert, 2023) reveal a bright, nearly symmetric emission around the Galactic Center (GC) and extending $\sim 6^{\circ}$$8^{\circ}$ on the sky. The inferred positron injection rate, $R_{e^{+}}\!\sim\!10^{43}\,{\rm s^{-1}}$ (Prantzos and others, 2011), together with the narrow extension of the signal, implies injection energies below a few–tens of MeV.
@@
While $\beta^{+}$ decays of 26Al, 44Ti and 56Ni can account for a substantial fraction of the disk component (Diehl and others, 2006; Prantzos and others, 2011), conventional astrophysical populations struggle to reproduce the bright, compact bulge emission (Siegert, 2023). Stellar templates such as the nuclear stellar bulge or a boxy bulge can be tuned to match the observed morphology (Siegert et al., 2021), which has been interpreted as support for a stellar origin. However, magnetohydrodynamic simulations of Milky Way–like galaxies show that dark matter (DM)–baryon interactions can drive the DM distribution into a configuration that closely traces the stellar bulge and boxy bulge structure (Muru et al., 2025). In such cases, template fits based solely on stellar tracers cannot reliably distinguish a purely stellar explanation from one in which DM contributes significantly.
@@
The DM interpretation of the 511 keV line has therefore received sustained attention (Boehm and others, 2004; Vincent et al., 2012; De la Torre Luque et al., 2025, 2024; De la Torre Luque; Pedro et al., 2025; Nguyen et al., 2025). Simple $s$-wave annihilation scenarios, in which MeV-scale DM directly produces $e^{+}e^{-}$ pairs, are tightly constrained by CMB and Big Bang Nucleosynthesis (Beacom and Yuksel, 2006; Sizun et al., 2007; Wilkinson et al., 2016; De la Torre Luque et al., 2025), disfavoring standard thermal MeV DM as the source of the positrons (however, see Ref. (Aghaie et al., 2025) for an interesting possibility which evades current constraints and is able to reproduce the 511 keV observations).
@@
A particularly promising alternative is the Excited Dark Matter (XDM) framework proposed in Finkbeiner and Weiner (2007). In this scenario, collisions between DM particles inelastically excite one particle to a slightly heavier state that decays back via a light mediator, producing an $e^{+}e^{-}$ pair. Because excitation requires a threshold in relative velocity, the positron production rate acquires an exponential dependence on the local velocity dispersion. As the halo velocity dispersion decreases with galactocentric radius, the upscattering rate is strongly suppressed outside the inner kiloparsec, yielding a bulge-peaked morphology that naturally declines toward the disk. Quantitative studies by Vincent et al. (2012) and Cappiello et al. (2023) showed that such XDM models can reproduce the INTEGRAL/SPI angular profile for TeV-scale DM masses, providing a markedly better fit than standard annihilating DM scenarios.
@@
In parallel, a reanalysis of archival COMPTEL data has uncovered a broad excess around 2–3 MeV, detected at $\sim 8\sigma$ significance and consistent with in-flight annihilation (IfA) of $\sim 2$ MeV positrons (Knödlseder et al., 2025). The spatial distribution of this MeV excess closely tracks the 511 keV emission, strongly suggesting a common source of low-energy positrons.
@@
A third anomaly arises in the Central Molecular Zone (CMZ), the dense neutral-gas complex within the central few hundred parsecs of the Galaxy. Multiple molecular tracers indicate an H2 ionization rate $\zeta_{\mathrm{H}_{2}}\gtrsim 10^{-15}\,\mathrm{s^{-1}}$ (Oka et al., 2005; Oka and Geballe, 2020), which is difficult to reconcile with standard CRs (cosmic rays): they are strongly attenuated in dense clouds and their predicted ionization rate falls short by at least two orders of magnitude (Ravikularaman et al., 2025; Huang et al., 2021). Moreover, ionization powered by a single central source such as Sgr A* or HESS J1745-290 is expected to peak sharply near the source and decline steeply with radius, whereas the data favour a much flatter radial profile across the CMZ (Dogiel et al., 2013; Ravikularaman et al., 2025). De la Torre Luque et al. (2025) showed that MeV DM annihilating into $e^{\pm}$ can efficiently ionize the CMZ and reproduce both the level and approximate flatness of the ionization profile, in a region of parameter space compatible with existing constraints and with positron injection rates comparable to those required for the 511 keV line.
@@
The remainder of this paper is organised as follows. Section II summarizes the XDM mechanism, introduces the key phenomenological parameters and describes our implementation of the excitation rate and positron source term. Section III presents the fit to the 511 keV line, and the prediction of the IfA signal and the CMZ ionization contribution. Section IV discusses the implications of our results and the prospects for testing this scenario with future MeV-band observations.
II Excited Dark Matter Framework
@@
We consider a minimal dark sector with two nearly degenerate states: a stable ground state $\chi$ and a slightly heavier excited state $\chi^{*}$, separated by a small mass splitting $\Delta m=m_{\chi^{*}}-m_{\chi}$. This setup, first proposed for the 511 keV line in Finkbeiner and Weiner (2007) and revisited in Vincent et al. (2012); Cappiello et al. (2023), provides a simple way for DM to produce low-energy positrons in the Milky Way.
Collisions between DM particles can inelastically upscatter one of them to the excited state, $\chi\chi\rightarrow\chi\chi^{*}$ provided their relative velocity exceeds a kinematic threshold. The excited state then decays $\chi^{*}\rightarrow\chi e^{+}e^{-}$, releasing an $e^{+}e^{-}$ pair with total energy $\sim\Delta m$. In contrast to standard annihilating DM, this mechanism converts a small fraction of the halo kinetic energy into electron–positron pairs, leaving the overall DM abundance and large-scale halo structure unchanged. Even if only a tiny fraction of particles are excited per Hubble time, the kinetic energy reservoir is sufficient to power the observed positron production rate in the Galactic bulge (Prantzos and others, 2011; Siegert, 2023).
Each upscattering event injects positrons with kinetic energy $E_{e^{+}}\simeq\Delta m/2$, so $\Delta m\sim\mathrm{few\penalty 10000\ MeV}$ naturally yields non-relativistic positrons consistent with the narrow 511 keV line. The strong velocity threshold induces a steep radial dependence in the excitation rate: because the velocity dispersion $\sigma_{v}(r)$ decreases with galactocentric radius, the upscattering rate is exponentially suppressed beyond the inner kiloparsec. This produces a sharply peaked, bulge-dominated morphology that matches the $\sim 6-8^{\circ}$ FWHM of the INTEGRAL/SPI signal (Weidenspointner and others, 2007; Siegert et al., 2019) without requiring ad hoc modifications of the DM density profile.
@@
as shown in Finkbeiner and Weiner (2007); Vincent et al. (2012); Cappiello et al. (2023). The exponential factor arises from integrating the inelastic cross section over the local Maxwell–Boltzmann velocity distribution and naturally suppresses the signal outside the Galactic bulge.
At the microscopic level, we parameterize the inelastic upscattering cross section as (Cappiello et al., 2023)
@@
where $\sigma_{\textrm{mr}}$ sets the normalization of the inelastic scattering cross section in the moderately relativistic limit. The threshold $v_{\mathrm{th}}$ introduces a sharp kinematic cutoff, confining excitations to the high-velocity tail of the halo distribution and imprinting the exponential suppression with radius seen in Eq. (2). This form approximates the more general expressions obtained in light-mediator models (Finkbeiner and Weiner, 2007; Pospelov and Ritz, 2007; Arkani-Hamed et al., 2009), while capturing the astrophysically relevant dependence needed to model the Galactic signal.
In what follows, we fold this threshold behaviour into a full computation of the velocity-averaged excitation rate using the local velocity dispersion $\sigma_{v}(r)$ derived from the Jeans equation (Section II). This quantity controls both the radial morphology of the 511 keV emission and the total positron injection rate relevant for the IfA and CMZ analyses.
@@
Velocity-averaged excitation rate. Following Cappiello et al. (2023), the velocity-averaged inelastic excitation cross section is obtained by integrating the microscopic cross section over the local DM velocity distribution,
@@
where $\sigma_{v}(r)$ and $v_{\mathrm{esc}}(r)$ are the local DM velocity dispersion and escape velocity, and $N_{\mathrm{esc}}$ normalizes the truncated Maxwellian distribution. The threshold velocity $v_{\mathrm{th}}$ is defined in Eq. (1). [^1] In Eq. (5) of Cappiello et al. (2023), a cross section $\langle\sigma v\rangle_{\rm mr}\simeq\sigma_{\textrm{mr}}v_{\rm eff}$ was introduced as a constant normalization with an effective halo velocity $v_{\rm eff}\sim 10^{-3}c$. Here we instead derive $\langle\sigma v\rangle(r)$ by explicitly integrating over the full velocity distribution, which is appropriate for a geometric cross section $\sigma_{\rm mr}$ and ensures a self-consistent radial dependence.
@@
The velocity dispersion $\sigma_{v}(r)$ and escape velocity $v_{\mathrm{esc}}(r)$ entering Eq. (5) are shown in Fig. 1. They are computed by solving the isotropic Jeans equation under spherical symmetry, following Vincent et al. (2012); Cappiello et al. (2023). The total gravitational potential includes contributions from DM, bulge, disk and Sgr A*. We adopt the same baryonic potential models and parameter choices as Vincent et al. (2012), which reproduce Milky Way rotation-curve data and yield $\sigma_{v}\sim 200$$250\penalty 10000\ \mathrm{km\,s^{-1}}$ and $v_{\mathrm{esc}}\sim 600$$700\penalty 10000\ \mathrm{km\,s^{-1}}$ in the inner kiloparsec.
Figure 1: DM velocity profiles as a function of radial distance from the GC for the velocity dispersion $\sigma_{v}$ (blue) and escape velocity $v_{\mathrm{esc}}$ (green) for NFW (solid) and cNFW (dashed) DM density distributions. The cNFW profile has inner slope $\gamma=1.2$.
@@
The angular distribution of 511 keV photons is obtained by integrating the local emissivity of diffuse positrons along the line of sight. Unlike previous morphology studies that assumed instantaneous annihilation at the production site (Vincent et al., 2012; Cappiello et al., 2023), we propagate positrons with non-zero kinetic energy using the DRAGON2 code, following De la Torre Luque et al. (2024). This includes diffusion, reacceleration and energy losses, and yields a steady-state positron distribution before annihilation.
@@
where $\frac{{d\phi_{e}}}{{dE}}(E_{\mathrm{th}})$ is the propagated positron flux at thermal energy $E_{\mathrm{th}}$, $\sigma^{\mathrm{ps}}$ is the charge-exchange cross section for positronium formation with hydrogen (Jean, P. and others, 2009), and $n_{e}$ is the ambient electron density. We take $n_{e}=1\penalty 10000\ \mathrm{cm^{-3}}$ in the Galactic plane and adopt an exponential decrease above and below the plane as in De la Torre Luque; Pedro et al. (2025), combining a Nakanishi warm-gas component (Nakanishi and Sofue, 2003) and a Ferriere hot-gas component (Ferriere, 1998).
@@
Relativistic positrons injected in the interstellar medium can annihilate with electrons before becoming thermal, producing a continuum of $\gamma$ rays up to the positron energy. IfA has long been recognized as a probe of MeV DM scenarios (Stecker, 1969; Beacom and Yuksel, 2006; Sizun et al., 2006), since high injection energies generate an IfA flux that can overshoot the observed diffuse MeV background (Strong et al., 1994; Bouchet et al., 2011).
As discussed in Sec. I, the reanalysis of COMPTEL data by Knödlseder et al. (2025) reveals a MeV excess consistent with IfA of $\sim 2$ MeV positrons sharing the 511 keV spatial template. In our XDM scenario, these positrons are produced in $\chi^{*}\to\chi e^{+}e^{-}$ decays with injection energy $E_{e^{+}}\simeq\Delta m/2$. For monoenergetic injection, the differential IfA $\gamma$-ray flux can be related to the 511 keV intensity (Beacom and Yuksel, 2006; Beacom et al., 2005):
@@
where $f_{\mathrm{Ps}}\simeq 0.97$ is the positronium fraction (Jean, P. and others, 2009; Bouchet et al., 2010) and $P^{\mathrm{IfA}}(E_{\gamma};E_{0},n_{e})$ is the photon spectrum per injected positron of initial energy $E_{0}=\Delta m/2$ in a medium of density $n_{e}$.
Here we compute the IfA signal following De la Torre Luque; Pedro et al. (2025), using the steady-state positron spectrum obtained from DRAGON2:
@@
where the first integral over $E^{\prime}$ accounts for the propagated positron energy distribution, bounded between $E_{\textrm{max}}$ (the maximum steady-state energy) and $m_{e}/2$ (Bartels et al., 2017). The factor $(1/N_{\rm pos})\,dN_{\rm pos}/dE^{\prime}$ gives the fraction of positrons at energy $E^{\prime}$ that contribute to the 511 keV emission. The hydrogen density $n_{\textrm{H}}$ sets the number of target electrons, $P_{E^{\prime}\to E}$ is the probability for a positron to produce a photon before cooling from $E^{\prime}$ to $E$ (Beacom and Yuksel, 2006), and $P=P_{E^{\prime}\to m_{e}}$ is the total probability to emit a photon before reaching thermal energies.
@@
As discussed in Sec. I, multiple molecular tracers indicate an H2 ionization rate in the CMZ of $\zeta_{\mathrm{H}_{2}}\gtrsim 10^{-15}\,\mathrm{s^{-1}}$ with an approximately flat radial profile (Oka et al., 2005; Oka and Geballe, 2020; Dogiel et al., 2013; Ravikularaman et al., 2025), which is difficult to explain with standard CRs alone (Huang et al., 2021; Ravikularaman et al., 2025). In light of the MeV–DM explanation proposed by De la Torre Luque et al. (2025), we now ask whether the same XDM setup that fits the 511 keV morphology and IfA signal can also contribute appreciably to the CMZ ionisation.
We follow De la Torre Luque et al. (2025) to compute the H2 ionization rate $\zeta_{\textrm{H}_{2}}$ caused by the $e^{\pm}$ pairs emitted in XDM de-excitations inside the CMZ. The rate is
@@
where the prefactor 2 accounts for ionization by both $e^{-}$ and $e^{+}$, $E_{\textrm{max}}=\Delta m/2$ is the injection energy, and $E_{\textrm{min}}$ is the threshold for ionization in H2 ($15.43$ eV (Padovani et al., 2009)). The cross section $\sigma(E)$ is the electron–H2 ionization cross section, calculated as in Eq. (7) of Padovani et al. (2009), and $\theta_{e}(E)$ is the number of secondary ionizations per primary ionization, from Eq. (2.23) of Krause et al. (2015). The flux $J(E,\textbf{x})$ is the propagated $e^{\pm}$ flux from DRAGON2, which includes the strong energy losses suffered by these particles in the CMZ. Typical propagation distances before thermalization are of order tens of parsecs, depending on the injection energy.